# Zero-Shot Clinical Trial Patient Matching with LLMs

### v1

@@ Abstract
Data and Code Availability
This paper uses the 2018 n2c2 Cohort Selection Challenge dataset (Stubbs et al., 2019). Our code is available here: https://github.com/som-shahlab/clinical_trial_patient_matching.
Institutional Review Board (IRB)
This was not determined to be human subject research, and thus did not require IRB approval.
@@ 1 Introduction
Figure 1: We explore zero-shot clinical trial patient matching with LLMs under two system designs: (a) We inject the patient’s entire set of notes into a prompt input into an Assessment LLM (e.g. GPT-4) for evaluation. (b) In our two-stage retrieval pipeline, we first query the top-$k$ most relevant chunks from the patient’s notes, then inject only those top-$k$ chunks into the prompt input into an Assessment LLM. Both paradigms are compared using the same prompting strategies, with the only distinction being the amount of patient information included in the prompt.
Identifying eligible patients for clinical trials is a key bottleneck in advancing new drugs to market (Woo, 2019). Roughly one third of clinical trials fails because it cannot enroll enough eligible patients, and recruitment costs eat up 32% of a trial’s budget (Deloitte, 2020).
@@
That is because identifying patients who are eligible for a trial is often highly manual and time-consuming (Brøgger-Mikkelsen et al., 2020). Each trial has a long list of eligibility criteria which must be exactly met by every patient who enrolls (FDA, 2020). To identify these patients, a trained clinical research coordinator manually reviews hundreds of patients’ electronic health records (EHRs) (Penberthy et al., 2012; Calaprice-Whitty et al., 2020). Over 80% of the relevant information is stored as unstructured text – e.g. progress notes, emails, radiology reports, and genetic tests (Kong, 2019) – making it difficult to automatically process (Leaman et al., 2015). As a result, eligibility screening for a Phase III cancer trial can take almost an hour per patient (Penberthy et al., 2012).
Traditional natural language processing (NLP) approaches have seen limited success with clinical text due to its highly idiosyncratic grammar and terminology (Leaman et al., 2015). Large language models (LLMs) represent a key inflection point in the capabilities of NLP systems (Brown et al., 2020), and thus offer a promising approach for accelerating patient recruitment.
In this paper, we aim to build upon recent work exploring the application of LLMs to patient matching (Jin et al., 2023; Wong et al., 2023; Nievas et al., 2023). In contrast to this prior work, our primary focus is on the efficiency of LLM-based approaches – namely, their cost in terms of dollars, money, and time. These considerations are essential if such systems are proven effective and aim to be deployed at scale.
@@
    We investigate the zero-shot (i.e. out-of-the-box) performance of LLMs on evaluating patient eligibility. Unlike prior work, this requires no training data or in-context examples. Our GPT-4-based system achieves state-of-the performance on Track 1 of the 2018 n2c2 cohort selection benchmark (Stubbs et al., 2019).
@@ 2.1 Large Language Models (LLMs)
Natural language processing (NLP) was revolutionized by the advent of deep learning models which could learn from huge amounts of unlabeled text, e.g. all of Wikipedia (Howard and Ruder, 2018). As these models increased in size and their datasets became even larger (e.g. all of the Internet), their capabilities continued to improve (Brown et al., 2020; Kaplan et al., 2020). Known as “large language models” (LLMs), they have surpassed humans on a number of benchmarks (Anil et al., 2023) and can now conduct diagnostic conversations with patients (Tu et al., 2024), translate clinical jargon into layman’s terms (Mirza et al., 2024), and perform chart review (Goel et al., 2023). Most impressively, LLMs have demonstrated the ability to tackle many of these tasks without ever explicitly being trained to solve them (Brown et al., 2020).
@@ 2.2 Clinical Trial Patient Matching
Initial work in automated trial matching focused on automatically transforming free text eligibility criteria into structured SQL queries, which could then be run over the structured data within patient EHRs. Systems such as EliXR (Weng et al., 2011) accomplished this via rule-based algorithms, while EliIE (Kang et al., 2017) and Criteria2Query (Yuan et al., 2019) used more advanced NLP methods such as named entity recognition, relation extraction, and negation detection. Most relevant to this paper, the RBC model from Oleynik et al. (2019) achieved the highest score on the 2018 n2c2 cohort selection benchmark. While well-suited for the specific trials for which they are developed, these systems are difficult to maintain and tend not to generalize (Gehrmann et al., 2017).
@@ 3 Methods
In the following, assume we are given a set of patients $\mathcal{P}$ and notes $\mathcal{N}$, where each patient $p\in\mathcal{P}$ has a set of notes $\mathcal{N}_{p}\in N$. We have a set $\mathcal{C}$ of trial criteria. There is some unknown function $E:\mathcal{C}\times\mathcal{P}\rightarrow\{0,1\}$ that maps each criterion and patient to a binary indicator of whether that patient meets (1) or does not meet (0) that criterion. Our goal is to calculate $E(c,p)$ for all patients $p$ and criteria $c$.
@@ 3.1 Models
For our assessment model $\mathbf{A}$, we consider the proprietary models GPT-3.5 Brown et al. (2020) and GPT-4 (Achiam et al., 2023), as well as the popular open source models Llama-2-70b-32k (Peng et al., 2023) and Mixtral-8x7B (Jiang et al., 2024). We run GPT-3.5 and GPT-4 via a secure Azure PHI-compliant instance. We run the open source models on a local PHI-compliant server with 4 Nvidia A100 GPUs using the vLLM (Kwon et al., 2023) and Outlines (Willard and Louf, 2023) packages for high-speed inference and enforcing valid JSON output.
Table 1: Zero-shot results on the 2018 n2c2 cohort selection challenge using the ACIN prompt strategy. We use the 32k context length version of all models, with the exception of GPT-3.5 (which is limited to 16k tokens).
@@
Table 2: Increasing the specificity of criteria definitions improves model accuracy.
@@
3.2 Zero-Shot Criteria Assessment
Given a set of notes $\mathcal{N}_{p}\subseteq\mathcal{N}$ for a patient $p$, we investigate four strategies for generating the prompt(s) that we feed into the assessment LLM: (1) All Criteria, All Notes (ACAN) in which all notes in $\mathcal{N}_{p}$ are merged into a single prompt and the model generates all $C=|\mathcal{C}|$ eligibility assessments at once. (2) All Criteria, Individual Notes (ACIN) in which the model is given $N=|\mathcal{N}_{p}|$ separate prompts (each containing one note), and for each note generates all $C$ eligibility assessments. (3) Individual Criteria, All Notes (ICAN) in which all $N$ notes are merged into a single prompt, but the model only assesses one criterion at a time, and thus must be re-prompted with the same merged set of notes $C$ times. (4) Individual Criteria, Individual Notes (ICIN) in which the model is given NCN*C separate prompts each containing one note and one criterion. More precisely, we thus have:
Where $\mathbin{\|}$ represents concatenation. Please see Appendix Figures 5 and 6 for the exact prompt templates used.
The output $o\in\mathcal{O}$ that we expect from the Assessment model $\mathbf{A}$ is a JSON string containing four elements:
  1. 1.
    Criterion: Specific criterion the model is assessing.
  2. 2.
    Medications: List of medications extracted from the note.
  3. 3.
    Rationale: Step-by-step reasoning as to how the model claims to have come to its decision.
  4. 4.
    Decision: Binary “MET” or “NOT MET” decision.
  5. 5.
    Confidence: ”Low”, ”Medium”, or ”High” confidence in prediction
We include the Medications element to helped the model recall specific drugs when assessing certain criteria, as prior literature demonstrated the ability of LLMs to perform this extraction (Agrawal et al., 2022). Rationale was included to both improve model performance via chain-of-thought-like reasoning (Wei et al., 2022) and provide interpretability for a human reviewer. For the Llama-2 model only, we removed Medications and Confidence from the prompt as we were unable to get it to generate valid completions.
3.3 Retrieval Pipeline
To initialize our pipeline, we loop through every patient and embed all of their notes using our embedding model $\mathbf{M}$. We also we embed each criterion.. Thus, we get a database of embeddings $\mathcal{D}$ such that $\mathcal{D}=\{\mathbf{M}(n)~\forall n\in\mathcal{N}\}\cup\{\mathbf{M}(c)~\forall c\in\mathcal{C}\}$
4 Evaluation
Table 3: Considering one criterion/note at a time improves performance. “Tokens” includes both prompt and completion tokens (i.e. inputs and outputs). “API Calls” is the total number of times the LLM was queried. “Cost” is based on OpenAI’s pricing as of January 25, 2024 (OpenAI, 2024)
@@ 4 Evaluation
Criteria Notes Prec. Rec.
Overall
Macro-F1
Overall
Micro-F1
Cost API Calls Tokens
@@
4.1 Dataset
For our dataset, we use the 2018 n2c2 Clinical Trial Cohort Selection challenge dataset, as it is one of the more well-studied public benchmarks for patient matching (Stubbs et al., 2019). The dataset has two parts: patients and eligibility criteria.
(1) Patients. There are 288 unique patients (subdivided into two splits: 202 train and 86 test). Each patient has 2-5 de-identified clinical notes, with an average of 2,711 words each. The patients are all diabetic, and were originally sourced from a combined database of patients from Mass General and Brigham and Women’s hospitals. Details on the cohort selection process are provided in Stubbs and Uzuner (2015) and Kumar et al. (2015).
(2) Eligibility criteria. There are 13 inclusion criteria which reflect commonly used eligibility criteria (Stubbs et al., 2019). These criteria, as well as their definitions, are listed in Appendix Table 5. Note that there are no exclusion criteria.
All of this information is provided as unstructured text. Thus, the only way to match patients is via natural language processing. Note that this benchmark simulates a synthetic trial – i.e. there is no actual clinical trial that enrolled these patients. To obtain labels, two medical experts annotated each patient as to whether the 13 inclusion criteria were “MET” or “NOT MET”. The average Cohen’s Kappa between the two annotators was 0.54 (Stubbs et al., 2019).
4.2 Metrics
We also measure several efficiency metrics: (1) total number of tokens (both prompt and completion) used by the LLM to generate its predictions; (2) total cost (in USD) of making predictions; (3) and total number of API calls made to the LLM (as a proxy for time taken).
4.3 Baselines
In addition to the LLMs we test, we consider two baselines. First, the best model from the original 2018 n2c2 competition, which achieved an Overall Micro-F1 of 0.91 (Oleynik et al., 2019). We refer to this model as “Prior SOTA.” It utilized hand-crafted rules, negation detection, family history detection, and regular expressions (Oleynik et al., 2019). We also consider simply predicting the class prevalence for each criterion, and refer to this as “Baseline”.
@@ 4.4 Interpretability
5 Results
Figure 2: Model performance increases as the number ($k$) of retrieved documents increases, but quickly plateaus with diminishing returns. We test $k\in\{1,3,5,10\}$. Each subfigure is a different prompting strategy. The y-axis is model performance (Macro/Micro-F1) and the x-axis is the total number of tokens processed by the model. Orange is GPT-4, blue is GPT-3.5, and the green line is the prior state-of-the-art. Stars represent each model’s best performance when feeding in all notes. The MiniLM embedding model is the dashed line, while BGE is the solid line.
@@ 5 Results
Figure 3: Clinician assessment of whether the rationales generated by GPT-4 were correct (green), partially correct (yellow), or incorrect (red), split by whether GPT-4’s corresponding eligibility decision was evaluated as correct (a) or incorrect (b) according to the n2c2 benchmark.
5.1 Zero-Shot Criteria Assessment
For our initial zero-shot evaluation, we feed the entire patient’s medical history into the LLM and have it predict all criteria at once per the ACIN prompting strategy. All of the models we test are able to fit each patient’s history into their context windows. We report our results in Table 1. Despite not being tuned for trial matching or provided any in-context examples, GPT-4 beats the state-of-the-art by a margin of +6 Macro-F1 and +2 Micro-F1 points. We provide criterion-level confusion matrices for this model in Appendix Figure 4.
The open source models perform worse than GPT-3.5, and GPT-3.5 performs significantly worse than GPT-4 (Bubeck et al., 2023). This aligns with prior evaluations of LLMs on EHR-based tasks (Fleming et al., 2023), and suggests that prior trial matching work based on GPT-3.5 has significantly underestimated the capabilities of LLMs for this task (Jin et al., 2023).
5.2 Prompt Engineering
The n2c2 benchmark does not come with a predefined set of prompts, but instead provides a set of sparse definitions for each criteria (see Appendix Table 5). As the low inter-rater agreement between clinicians suggests, these definitions did not fully capture how the dataset was labeled. Thus, to design our prompts, we randomly sampled three patients who met and did not meet each eligibility criteria from the training dataset, and used this corpus to manually refine the definition of each criterion. We provide our “improved” criteria in Appendix Table 6, and demonstrate in Table 2 that increasing the specificity of criteria is an essential first step in using an LLM. In a real-world setting, this step could be avoided by directly using the detailed protocols provided by trial sponsors (DNDi, 2018; Limited, 2017).
We ablate the 4 prompting strategies detailed in Section 3.2. As shown in Table 3, the ACAN strategy performs worst, while the rest of the strategies appear to be relatively equivalent in performance. When cost is considered, however, the ACIN strategy becomes the clear winner – the ACIN strategy achieves the best Micro-F1 at almost an order of magnitude lower cost, API calls, and tokens used than the other prompting strategies. ACIN passes each note through the LLM once and has it evaluate all criteria simultaneously, whereas the strategies relying on “Individual Criteria” pass each note through the LLM 13 times (once for each criterion). Thus, careful prompting can result in meaningful cost savings with virtually no drop in accuracy.
5.3 Retrieval Pipeline
In Figure 2, we plot the performance of the model on the y-axis and the amount of tokens used on the x-axis. In these plots, the orange line is GPT-4, the blue line is GPT-3.5, the green line is the prior state-of-the-art, and the star is the non-retrieval-based implementation of each model. With our best embedding model, MiniLM, we are able to surpass the prior state-of-the-art on Macro-F1 using roughly one-third and one-half as many tokens as needed in the vanilla ICAN and ACAN strategies, respectively. While both methods approach the state-of-the-art on Micro-F1, neither retrieval-based method is able to fully achieve it without using the full note. There appears to be an unresolved gap in performance when relying on a retrieval-based system to curate patient information. Additionally, the cost benefits of the retrieval pipeline largely depend on the prompting strategy used – since the ACIN strategy feeds each chunk separately through the model, once $k$ exceeds the number of total notes in a patient’s EHR, then the cost of using the retrieval pipeline ends up exceeding the cost of simply using full notes without retrieval.
5.4 Interpretability
@@
We sample 468 rationales generated by GPT-4 using the ICAN prompting strategy, and have two clinicians evaluate their veracity. We use the ICAN version of the system in order to cleanly disambiguate each individual criterion’s prediction. Each rationale was evaluated on a 3-part scale: Correct, Partially Correct, and Incorrect, based on how accurately it aligned to the relevant patient’s EHR.
@@
The results in Figure 3 show that GPT-4 is able to provide legitimate rationales for most its decisions. When GPT-4 makes a correct eligibility decision, 89% of its rationales were judged as fully correct, 8% as partially correct, and 3% as incorrect. When GPT-4 made an incorrect eligibility decision, its rationales were split 67/8/25% between correct/partially correct/incorrect. This indicates that either the “ground truth” label of the n2c2 benchmark is incorrect, or that there is legitimate variation in clinical judgement as to whether the patient meets a criterion. This nuance is something that an LLM can better capture through its natural language generation abilities, whereas traditional rule-based and deep learning models cannot.

### v2

@@ 1 Introduction
Figure 1: Zero-Shot system design
We explore zero-shot clinical trial patient matching with LLMs under two system designs: (a) We inject the patient’s notes into a prompt that is fed to an Assessment LLM (e.g. GPT-4) for evaluation. The ACAN prompting strategy is depicted here. (b) In our two-stage retrieval pipeline, we first query the top-$k$ most relevant chunks from the patient’s notes, then inject only those top-$k$ chunks into the prompt fed to the Assessment LLM. Again, the ACAN prompting strategy is depicted here.
Figure 1: Zero-Shot system design
Identifying eligible patients for clinical trials is a key bottleneck in advancing new drugs to market (Woo, 2019). One third of clinical trials fail because they cannot enroll enough patients, and recruitment costs an average of 32% of a trial’s budget (Deloitte, 2020).
@@
That is because identifying patients who are eligible for a trial is often highly manual and time-consuming (Brøgger-Mikkelsen et al., 2020). Each trial has a long list of eligibility criteria which must be exactly met by every patient who enrolls (FDA, 2020). To identify these patients, a trained clinical research coordinator will manually review hundreds of patients’ electronic health records (EHRs) (Penberthy et al., 2012; Calaprice-Whitty et al., 2020). Over 80% of the relevant information is stored as unstructured text – e.g. progress notes, emails, radiology reports, and genetic tests (Kong, 2019) – making it difficult to automatically process (Leaman et al., 2015). As a result, eligibility screening for a Phase III cancer trial can take almost an hour per patient (Penberthy et al., 2012).
Traditional natural language processing (NLP) approaches have seen limited success with clinical text due to its idiosyncratic grammar and terminology (Leaman et al., 2015). Large language models (LLMs) represent a key inflection point in NLP capabilities (Brown et al., 2020), and thus offer a promising approach for accelerating patient recruitment.
In this paper, we aim to build upon recent work exploring the application of LLMs to patient matching (Jin et al., 2023; Wong et al., 2023; Nievas et al., 2023). In contrast to this prior work, our primary focus is on the efficiency of LLM-based approaches – namely, their cost in terms of dollars, money, and time. These considerations are essential if such systems aim to be deployed at scale given the limited resources of most health systems.
@@
    We investigate the zero-shot (i.e. out-of-the-box) performance of LLMs on evaluating patient eligibility. The zero-shot approach that we study offers several advantages over alternative such as fine-tuning or few-shot prompting, as the former requires no labeled data and would thus allow our system to more readily scale to arbitrary trials. It is also significantly cheaper than fine-tuning (which require purchasing compute) or few-shot prompting (which inflates token counts by a factor of over 2x in experiments). Our zero-shot GPT-4-based system achieves state-of-the performance on Track 1 of the 2018 n2c2 cohort selection benchmark, which is the largest and most realistic publicly available clinical trial patient matching benchmark (Stubbs et al., 2019).
@@ 2.1 Large Language Models (LLMs)
Natural language processing (NLP) was revolutionized by the advent of deep learning models which could learn from huge amounts of unlabeled text, e.g. all of Wikipedia (Howard and Ruder, 2018). As these models increased in size and their datasets became even larger (e.g. all of the Internet), their capabilities continued to improve (Brown et al., 2020; Kaplan et al., 2020). Known as “large language models” (LLMs), they have surpassed humans on a number of benchmarks (Anil et al., 2023) and can now conduct diagnostic conversations with patients (Tu et al., 2024), translate clinical jargon into layman’s terms (Mirza et al., 2024), and perform chart review (Goel et al., 2023). Most impressively, LLMs have tackled many of these tasks without ever explicitly being trained to solve them (Brown et al., 2020).
@@ 2.2 Clinical Trial Patient Matching
Initial work in automated trial matching focused on automatically transforming free text eligibility criteria into structured SQL queries, which could then be run over the structured data within patient EHRs. Systems such as EliXR (Weng et al., 2011) accomplished this via rule-based algorithms, while EliIE (Kang et al., 2017) and Criteria2Query (Yuan et al., 2019) used more advanced NLP methods such as named entity recognition, relation extraction, and negation detection. Most relevant to this paper, the RBC model from (Oleynik et al., 2019) achieved the highest score on the 2018 n2c2 cohort selection benchmark. While well-suited for the specific trials for which they are developed, these systems are difficult to maintain and tend not to generalize (Gehrmann et al., 2017).
@@ 3 Methods
In the following, assume we are given a set of patients $\mathcal{P}$ and notes $\mathcal{N}$, where each patient $p\in\mathcal{P}$ has a set of notes $\mathcal{N}_{p}\subseteq\mathcal{N}$. We have a set $\mathcal{C}$ of trial criteria. There is some unknown function $E:\mathcal{C}\times\mathcal{P}\rightarrow\{0,1\}$ that maps each criterion and patient to a binary indicator of whether that patient meets (1) or does not meet (0) that criterion. Our goal is to calculate $E(c,p)$ for all patients $p$ and criteria $c$.
@@ 3.1 Models
For our assessment model $\mathbf{A}$, we consider the proprietary models GPT-3.5 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), as well as the popular open source models Llama-2-70b-32k (Peng et al., 2023) and Mixtral-8x7B (Jiang et al., 2024). We run GPT-3.5 and GPT-4 via a secure Azure PHI-compliant instance. We run the open source models on a local PHI-compliant server with 4 Nvidia A100 GPUs using the vLLM (Kwon et al., 2023) and Outlines (Willard and Louf, 2023) packages for inference and enforcing valid JSON output.
3.2 Zero-Shot Criteria Assessment
Given a set of notes $\mathcal{N}_{p}\subseteq\mathcal{N}$ for a patient $p$, we investigate four strategies for generating the prompt(s) that we feed into the assessment LLM: (1) All Criteria, All Notes (ACAN) in which all notes in $\mathcal{N}_{p}$ are merged into a single prompt and the model generates all $C=|\mathcal{C}|$ eligibility assessments at once. (2) All Criteria, Individual Notes (ACIN) in which the model is given $N=|\mathcal{N}_{p}|$ separate prompts (each containing one note), and for each note generates all $C$ eligibility assessments. (3) Individual Criteria, All Notes (ICAN) in which all $N$ notes are merged into a single prompt, but the model only assesses one criterion at a time, and thus must be re-prompted with the same merged set of notes $C$ times. (4) Individual Criteria, Individual Notes (ICIN) in which the model is given $N*C$ separate prompts each containing one note and one criterion. More precisely, we thus have:
[…]
The output $o\in\mathcal{O}$ that we expect from the Assessment model $\mathbf{A}$ is a JSON string containing five elements:
  1. 1.
    Criterion: The specific criterion being assessed.
  2. 2.
    Medications: The names of all current medications and supplements that the patient is taking.
  3. 3.
    Rationale: Step-by-step reasoning as to why the patient does or does not meet the criterion.
  4. 4.
    Decision: Output “MET” if the patient meets the criterion, or it can be inferred that they meet the criterion with common sense. Output “NOT MET” if the patient does not, or it is impossible to assess given the provided information.
  5. 5.
    Confidence: “Low”, “Medium”, or “High” confidence in prediction.
We include the Medications element to help the model recall specific drugs when assessing certain criteria, as prior literature demonstrated the ability of LLMs to perform this extraction (Agrawal et al., 2022). Rationale was included to provide interpretability for a human reviewer. We ablate this in Section 5.1 and find its inclusion has a slightly negative impact on accuracy. For the Llama-2 model only, we removed Medications and Confidence from the prompt as we were unable to get it to generate valid completions.
3.3 Retrieval Pipeline
To initialize our pipeline, we loop through every patient and embed all of their notes using our embedding model $\mathbf{M}$. We also embed each criterion. Thus, we get a database of embeddings $\mathcal{D}$ such that $\mathcal{D}=\{\mathbf{M}(n)~{}\forall n\in\mathcal{N}\}\cup\{\mathbf{M}(c)~{}\forall c\in\mathcal{C}\}$.
4 Evaluation
4.1 Dataset
We chose the 2018 n2c2 Clinical Trial Cohort Selection challenge dataset for evaluation as it is the largest and most realistic publicly available benchmark for clinical trial patient matching (Stubbs et al., 2019). Please see Appendix Section B for a more detailed comparison to other benchmarks. This dataset has two parts: patients and eligibility criteria.
(1) Patients. There are 288 unique patients (subdivided into two splits: 202 train and 86 test). Each patient has 2-5 de-identified clinical notes, with an average of 2,711 words each. The patients are all diabetic, and were originally sourced from a combined database of patients from Mass General and Brigham and Women’s hospitals. Details on the cohort selection process are provided in (Stubbs and Uzuner, 2015) and (Kumar et al., 2015).
(2) Eligibility criteria. There are 13 inclusion criteria which reflect commonly used eligibility criteria (Stubbs et al., 2019). These criteria, as well as their definitions, are listed in Appendix Table LABEL:tab:criteria_definition. Note that there are no exclusion criteria.
All of this information is provided as unstructured text. Thus, the only way to match patients is via natural language processing. Note that this benchmark simulates a synthetic trial – i.e. no actual clinical trial enrolled these patients. To obtain labels, two medical experts annotated each patient as to whether the 13 inclusion criteria were “MET” or “NOT MET”. The Cohen’s Kappa between the two annotators was 0.54 (Stubbs et al., 2019).
4.2 Metrics
We also measure several efficiency metrics: (1) total number of tokens (both prompt and completion) used by the LLM; (2) total cost (in USD) of making predictions; (3) and total number of calls made to the LLM (as a proxy for time).
4.3 Baselines
In addition to LLMs, we consider two baselines. First, the best model from the original 2018 n2c2 competition, which achieved an Overall Micro-F1 of 0.91 (Oleynik et al., 2019). We refer to this model as “Prior SOTA.” It utilized hand-crafted rules, negation detection, family history detection, and regular expressions (Oleynik et al., 2019). We also consider predicting the class prevalence for each criterion, and refer to this as “Baseline”.
4.4 Interpretability
5 Results
5.1 Zero-Shot Criteria Assessment
For our initial zero-shot evaluation, we follow the ACIN prompting strategy by feeding each note separately into the LLM and have it predict all criteria at once. All of the models we test are able to fit each patient’s history into their context windows. We report our results in Table LABEL:tab:zero_shot. Despite not being tuned for trial matching or provided any in-context examples, GPT-4 beats the state-of-the-art by a margin of +6 Macro-F1 and +2 Micro-F1 points. We provide criterion-level confusion matrices for this model in Appendix Figure LABEL:fig:confusion_matrices. The open source models perform worse than GPT-3.5, and GPT-3.5 performs significantly worse than GPT-4 (Bubeck et al., 2023). This aligns with prior evaluations of LLMs on EHR-based tasks (Fleming et al., 2023), and suggests that prior trial matching work based on GPT-3.5 has significantly underestimated the capabilities of LLMs for this task (Jin et al., 2023).
@@ 5.1 Zero-Shot Criteria Assessment
Table 1: Zero-shot results on n2c2 2018 challenge
While we have the model output a Rationale with each prediction for interpretability, we find that the inclusion of this field has a slightly negative impact on overall accuracy (results in Appendix Table LABEL:tab:ablation_rationale). As a brief initial comparison with few-shot prompting, we also evaluate the impact of including one example in the prompt. Surprisingly, we find in Appendix Table LABEL:tab:few_shot_1 that our simple 1-shot prompting approach performs worse than zero-shot prompting. We hypothesize that this is due to the substantially higher number of tokens included in the 1-shot prompts (roughly 2.45x more for the ACIN strategy), most of which are noise given the nature of clinical text. This makes it more difficult for GPT-4 to parse the relevant portions of the clinical note it is actually tasked with evaluating. Better prompting and fine-tuning strategies could certainly increase performance; however, given the focus of this paper is on the data/cost efficiency of LLMs, we leave such exploration to future work.
5.2 Prompt Engineering
The n2c2 benchmark does not come with a predefined set of prompts, but instead provides a brief definition for each criteria (see Appendix Table LABEL:tab:criteria_definition). As the low inter-rater agreement between clinicians suggests, these definitions did not fully capture how the dataset was labeled. Thus, to design our prompts we randomly sampled three patients who met and did not meet each criteria from the training dataset, and used this corpus to manually refine the definition of each criterion. We provide our “improved” criteria in Appendix Table LABEL:tab:criteria_improved, and demonstrate in Table LABEL:tab:prompt_eng that increasing the specificity of criteria is an essential first step in using an LLM. In the real-world, this step could be avoided by directly using the detailed protocols provided by trial sponsors (DNDi, 2018; Limited, 2017).
@@ 5.2 Prompt Engineering
Table 2: Increasing prompt specificity improves performance
We ablate the 4 prompting strategies detailed in Section 3.2. As shown in Table LABEL:tab:prompt_strat, the ACAN strategy performs worst, while the rest of the strategies appear to be relatively equivalent in performance. When cost is considered, however, the ACIN strategy becomes the clear winner – the ACIN strategy achieves the best Micro-F1 at almost an order of magnitude lower cost, API calls, and tokens used than the other prompting strategies. ACIN passes each note through the LLM once and has it evaluate all criteria simultaneously, whereas the strategies relying on “Individual Criteria” pass each note through the LLM 13 times (once for each criterion). Thus, careful prompting can result in meaningful cost savings with virtually no drop in accuracy.
@@
Criteria Notes Prec. Rec. Overall Macro-F1 Overall Micro-F1 Cost API Calls Tokens
@@
Table 3: Cost and data efficiency of prompting strategies
5.3 Retrieval Pipeline
In Figure LABEL:fig:retrieval, we plot the performance of the model on the y-axis and the amount of tokens used on the x-axis. In these plots, the orange line is GPT-4, the blue line is GPT-3.5, the green line is the prior state-of-the-art, and the star is the non-retrieval-based implementation of each model. With our best embedding model, MiniLM, we are able to surpass the prior state-of-the-art on Macro-F1 using roughly one-third and one-half as many tokens as needed in the vanilla ICAN and ACAN strategies, respectively. While both methods approach the state-of-the-art on Micro-F1, neither retrieval-based method is able to fully achieve it without using the full note. There appears to be an unresolved gap in performance when relying on a retrieval-based system to curate patient information. Additionally, the cost benefits of the retrieval pipeline largely depend on the prompting strategy used – as shown in Appendix Figure LABEL:fig:retrieval_acin since the ACIN strategy feeds each chunk separately through the model, once $k$ exceeds the number of total notes in a patient’s EHR, then the cost of using the retrieval pipeline ends up exceeding the cost of simply using full notes without retrieval. That is because each prompt contains a high number of non-note tokens (e.g. the instructions and criteria definitions), so splitting a note that could have been fit into one prompt into chunks spread across multiple separate prompts will increase the overall number of tokens used.
Figure 2: Two-stage retrieval pipeline performance
@@ 5.4 Interpretability
Figure 3: Clinician assessment of LLM-generated rationales
@@
We sample 486 rationales generated by GPT-4 using the ICAN prompting strategy, and have two clinicians evaluate their veracity. We use the ICAN version of the system in order to cleanly disambiguate each individual criterion’s prediction. Each rationale was evaluated on a 3-part scale: Correct, Partially Correct, and Incorrect, based on how accurately it aligned to the relevant patient’s EHR.
@@
The results in Figure LABEL:fig:clinican_rationale show that GPT-4 is able to provide legitimate rationales for most its decisions. When GPT-4 makes a correct eligibility decision, 89% of its rationales were judged as fully correct, 8% as partially correct, and 3% as incorrect. When GPT-4 made an incorrect eligibility decision, its rationales were split 67/8/25% between correct/partially correct/incorrect. This indicates that either the “ground truth” label of the n2c2 benchmark is incorrect, or that there is legitimate variation in clinical judgement as to whether the patient meets a criterion. This nuance is something that an LLM can capture through natural language in contrast to traditional rule-based and deep learning models.